OpenStudy (anonymous):

8^x+3=16^x−1 solve the exponential equation

OpenStudy (anonymous):

\[8^{(x+3)} = 16^{(x-1)}\] Let the base equal. \[2^{(3x+3)} = 2^{4(x-1)}\] Then, just set the exponents equal to each other. \[3(x+3) = 4(x-1)\] \[3x+9 = 4x - 4\] \[4x-3x = 9+4\] \[x = 13\]

OpenStudy (anonymous):

thanks dear

OpenStudy (alfie):

I was trying to solve the one he wrote and I was going INSANE... lol...!

OpenStudy (anonymous):

Fix :: \[2^{3(x+3)} = 2^{4(x-1)}\] **

OpenStudy (alfie):

\[8^x -16^x +4\] I was searching solutions for this one :D...!

OpenStudy (anonymous):

hahahaha @alfie

OpenStudy (anonymous):

i guess i posted it the wrong way?!

OpenStudy (anonymous):

It should have been written as: 8^(x+3)=16^(x−1) It is clearer.

OpenStudy (anonymous):

ok thanks